Determine the location of all critical points and determine their nature for the function. f(x)=ln(x^2 + 1) - x


1
Expert's answer
2021-06-10T06:55:57-0400

Here "f(x)=ln(x^2+1)-x"

We need to find location of all critical points & their nature for the function. For that we need to put "f'(x)=0"


Here "f'(x)=\frac {d} {dx} (ln(x^2+1)-x)"


"=\frac {2x} {x^2+1}" "-1" "\because \frac {d} {dx} (ln (x))=\frac 1 x" & "\frac {d} {dx} (x^n)=nx^{n-1}"


To find critical points we will put "f'(x)=0"

So "\frac {2x} {x^2+1} -1=0"


"\implies \frac {2x} {x^2+1} =1"

"\implies 2x=x^2+1"

"\implies x^2-2x+1=0"

"\implies (x-1)^2=0"

"\implies x=1"

So at "x=1" we get critical point & when "x=1" , "f(1)=ln(1+1)-1=ln2-1"


So critical point is "(1,ln2 -1)"


Now "f''(x)=\frac {d} {dx} (\frac {2x} {x^2+1} -1)=\frac {2-2x^2} {(x^2+1)^2}"


"\because \frac {d} {dx} (\frac u v)=\frac {(v.u'-u.v')} {v^2}" & "\frac {d} {dx} (constant)=0"


Now "f''(1)=\frac {(2-2)} {(2)^2} =\frac 0 4 =0"

As "f''(1)=0" So we will get inflection point at x=1. That is there is no local Maxima and local minima . (Ans)